### v1

# Vosh: Voxel-Mesh Hybrid Representation for Real-Time View Synthesis

Abstract.
The neural radiance field (NeRF) has emerged as a prominent methodology for synthesizing realistic images of novel views. While neural radiance representations based on voxels or mesh individually offer distinct advantages, excelling in either rendering quality or speed, each has limitations in the other aspect. In response, we propose a pioneering hybrid representation named Vosh, seamlessly combining both voxel and mesh components in hybrid rendering for view synthesis. Vosh is meticulously crafted by optimizing the voxel grid of NeRF, strategically with selected voxels replaced by mesh. Therefore, it excels in fast rendering scenes with simple geometry and textures through its mesh component, while simultaneously enabling high-quality rendering in intricate regions by leveraging voxel component. The flexibility of Vosh is showcased through the ability to adjust hybrid ratios, providing users the ability to control the balance between rendering quality and speed based on flexible usage. Experimental results demonstrates that our method achieves commendable trade-off between rendering quality and speed, and notably has real-time performance on mobile devices.
Neural Radiance Field, View Synthesis, Real-Time Rendering.
journal: TOGcopyright: noneccs: Computing methodologies Reconstructionccs: Computing methodologies Neural networksccs: Computing methodologies Volumetric models
Figure 1. Left: The hybrid representation Vosh, combines both voxel component (highlighted with pseudo colors) and mesh component (indicated by gray triangles) in hybrid rendering. Right: The proposed method facilitates real-time view synthesis on various mobile devices, e.g., laptops and mobile phones.
1. Introduction
The novel view synthesis based on the Neural Radiance Field (NeRF) leverages neural networks to model three-dimensional scenes and generates images from specified viewpoints (Mildenhall et al., 2021). In general, NeRF-based approaches demonstrate proficiency in producing high-quality rendering for novel views with sparse inputs, effectively handling materials with specular and translucent properties. This versatility positions them for diverse applications in autonomous navigation (Kwon et al., 2023), virtual reality/augmented reality (Li et al., 2022a), and beyond (Sun et al., 2022b; Jiang et al., 2023). Despite these merits, current methods face challenges in achieving a balance between high-quality rendering and real-time rendering, particularly when deployed on mobile devices.
The rendering speed of neural radiance fields is significantly influenced by the chosen representation, with notable implications for sampling requirements and query costs (Yu et al., 2021). Typically, the original NeRF uses deep multilayer perceptrons (MLPs) to model neural radiance fields, relying on dense sampling for querying to approximate integrals in volume rendering. Unfortunately, this approach typically causes excessively long rendering times. A substantial amount of effort has explored alternatives such as implicit voxel representations (Sun et al., 2022a; Hedman et al., 2021; Liu et al., 2020; Fridovich-Keil et al., 2022; Müller et al., 2022; Reiser et al., 2021) and explicit mesh representations (Chen et al., 2023; Tang et al., 2023; Rakotosaona et al., 2023; Yariv et al., 2023), aiming to exclude irrelevant regions for computation. These methods reduce the number of sampling points and employ smaller MLPs to expedite rendering. However, even though approaches based on implicit voxels can achieve high-quality rendering in complex scenes, the dense sampling in non-empty regions is prohibitively expensive.
Methods based on mesh representation have demonstrated significantly accelerated rendering speeds, enabling real-time rendering on common mobile devices. However, the challenge lies in the reconstruction of scene-level mesh structure, particularly in areas with intricate geometry and textures (see Figure 1), such as thin structures or undetermined regions like backgrounds (Reiser et al., 2023). It is observed that meshes can match the rendering quality akin to voxels in regions with simple geometry and textures, as well as offering highly efficient rendering. Therefore, a promising solution involves employing hybrid representations for reconstruction towards both high-quality rendering and fast speed. Existing works on hybrid representations (Wang et al., 2023; Guo et al., 2023), focus primarily on object or proximal region reconstruction, thus limiting flexibility to represent complex scenes and control over different parts of the hybrid representation. Alternatively, we leverage the strengths of voxels and mesh, to design a hybrid representation towards a better balance between quality and speed, especially enabling high-quality rendering on mobile devices (see Figure 1).
In summary, the main contribution of this paper is a new hybrid representation namely Vosh (VOxel-meSH), by combing voxels and mesh for NeRF-based real-time view synthesis. It also allows for a dynamic control of the balance between voxel and mesh components by voxel adjustment. Such flexibility results in a versatile spectrum of NeRF representations, catering to the varied computational capabilities that facilitate real-time view synthesis across different devices. Experimental results exhibit the super performance of our method compared to state-of-the-art methods.
Figure 2. An overview of the proposed methodology. The training phase starts from grid training for obtaining initial voxels. Then, a portion of voxels becomes mesh in the voxels to mesh conversion. Subsequently, the combination of voxels and mesh are optimized through hybrid rendering and voxels adjustment to obtain the final hybrid representation Vosh. The inference phase realizes real-time hybrid rendering with Vosh even on mobile phones.
2. Related Work
There are numerous published approaches to view synthesis and NeRF, and we refer interested readers to (Tewari et al., 2022; Gao et al., 2022) for a comprehensive overview. In this context, we focus on the real-time view synthesis and NeRF representations most related to our method.
2.1. Real-time view synthesis
NeRF (Mildenhall et al., 2021) as a well-known technology for view synthesis, has been the focus of numerous endeavors aimed at enhancing its rendering speed while maintaining high-quality rendering. SNeRG (Hedman et al., 2021) achieves fast rendering without CUDA, using sparse voxel representations and a compact neural network for deferred rendering. Plenoxels (Fridovich-Keil et al., 2022) and PlenOctrees (Yu et al., 2021) expedite ray marching with sparse voxel representation, refining voxels near object surfaces. MERF (Reiser et al., 2023) enhances rendering speed and spatial occupancy by replacing pure voxel grids with a low-resolution voxel grid and a high-resolution triplane projection representation (Chan et al., 2022). These methods mostly excel on high-end GPUs, whereas real-time rendering on mobile devices still remains a challenge.
MobileNeRF (Chen et al., 2023) achieves fast rendering by constraining NeRF to triangles and using rasterization. NeRF2Mesh (Tang et al., 2023) uses NeRF to initialize geometric structures, iteratively refining mesh attributes to improve rendering quality. BakedSDF (Yariv et al., 2023) optimizes a blended neural signed distance function representation, enabling real-time rendering of complex scenes on a high-quality mesh. While mesh-based methods offer fast rendering, they may suffer quality loss compared to voxel-based methods.
3D Gaussian Splatting (Kerbl et al., 2023) creates a 3D Gaussian representation around each point starting from a point cloud model, utilizing the Splatting method for real-time rendering of high-resolution scenes. But it suffers elongated and popping artifacts, and high memory usage for deployment on mobile phones. In contrast, our hybrid representation combines the strengths of voxel and mesh representations, allowing high-quality rendering of complex scenes on mobile devices.
2.2. Hybrid Scene Representation
Combining explicit mesh representation with implicit fields introduces functionalities such as NeRF editing and rendering acceleration. NeuMesh (Yang et al., 2022) encodes a neural implicit field with disentangled geometry and texture on mesh vertices, enabling a range of neural rendering-based editing operations. EyeNeRF (Li et al., 2022b) integrates an explicit surface with an implicit volume representation to model the human eye. Mesh-Aware-RF (Qiao et al., 2023) embeds a known polygonal mesh into NeRF, employing ray tracing and ray marching alternately to update radiance along rays.
The hybrid representation like Adaptive Shells (Wang et al., 2023) and VMesh (Guo et al., 2023) has been used in view synthesis. Adaptive Shells confines volume rendering to a narrow band around the object for effective NeRF rendering. VMesh, closely related to our work, uses a textured mesh and auxiliary sparse voxels but differs by employing SDF field and volume density field to simultaneously learn the entire scene. However, VMesh relies on gradient descent optimization that lacks controllability and flexibility. Indeed, VMesh mainly focuses on rendering individual objects or bounded scenes, while the proposed method in this paper can handle more scenes, such as unbounded outdoor scenes.
3. Method Overview
We start with an overview of our methodology to establish a context for the core algorithms described in the next section. As depicted in Figure 2, our method contains two key phases: the training phase for constructing a hybrid representation and the inference phase for view synthesis rendering. In the training phase, a set of multi-view images is employed to train a voxel-mesh hybrid representation, and the associated assets are exported. In the inference phase, our method achieves real-time rendering based on the baked hybrid representation assets.
Concretely, the training phase consists of three stages: grid training, voxels to mesh conversion and optimization. In the grid training stage, we train a voxel grid model through volume rendering. Then, leveraging the trained voxel grid, we extract an explicit surface and train the surface using differentiable surface rendering for both appearance and geometry optimization. This converts the initial voxel grid into a combination of voxel and mesh components. Finally, the voxels and mesh are further optimized through a training process involving hybrid rendering and voxels adjustment to obtain the desired hybrid representation. Once training is complete, we export assets containing sparse voxels and textured meshes. Subsequently, real-time view synthesis on mobile devices is implemented using WebGL. Next, we elaborate the algorithm details in the training and inference phases.
4. Hybrid Representation construction
The construction of a hybrid representation in the training phase commences with the initialization of a volume comprising solely of voxels. This volume is then subject to a conversion process wherein a portion of the voxels is replaced with an explicit mesh according to rendering quality on par with the original voxel representation.
4.1. Grid training for initial voxels
Initially, the entire scene is reconstructed using a high-resolution voxel grid. It retains sufficient geometric and appearance information essential for subsequent mesh extraction and hybrid representation. Here, we employ SNeRG++ (an improved version of SNeRG (Hedman et al., 2021) in MERF (Reiser et al., 2023)) as the benchmark for our voxel grid, due to its capability to effectively model unbounded scenes and generate real-time rendering assets.
Concretely, SNeRG++ is built upon the volume rendering principles of vanilla NeRF, which involves sampling along the direction of a ray originating from the ray origin $o$ along the direction $d$ to obtain sample points $\mathrm{\mathbf{x}}_{i}=o+t_{i}\mathrm{\mathbf{d}}$. Based on the spatial position of these sample points, it uses the neural radiance field to calculate volumetric density $\sigma_{i}$, resulting in rendering weights $w_{i}$:
(1) $w_{i}=\alpha_{i}T_{i},\;\;\;\;\;\;\;\;T_{i}=\prod^{i-1}_{j=1}(1-\alpha_{j}),\;\;\;\;\;\;\;\;\alpha_{i}=1-e^{-\sigma_{i}\delta_{i}},$
where $T_{i}$, $\alpha_{i}$ and $\delta_{i}$ represent the transmittance, opacity and spacing of sampling point $i$, respectively. SNeRG++ adopts the concept of deferred rendering, decomposing the appearance information of the neural radiance field into diffuse color $\mathrm{\mathbf{c}}_{d,i}$ and view-dependent features $f_{i}$. By cumulatively sampling the appearance information along the ray, as shown in Figure 4a, we obtain the diffuse color $C_{d}$ and view-dependent features $F$ of the ray as
(2) $\mathrm{\mathbf{C}}_{d}=\sum_{i}w_{i}\mathrm{\mathbf{c}}_{d,i},\;\;\;\;\;\;\;\;\mathrm{\mathbf{F}}=\sum_{i}w_{i}\mathrm{\mathbf{f}}_{i}.$
Finally, the view-dependent color is calculated by inputting the diffuse color $\mathrm{\mathbf{C}}_{d}$, view-dependent features $F$, and the ray direction $d$ into a miniature MLP. The result is then added to the diffuse color $\mathrm{\mathbf{C}}_{d}$ to obtain the final color $C$ of the ray.
Besides, in order to enhance the modeling of regions situated far from the center within unbounded scenes and facilitate efficient ray marching calculations, SNeRG++ incorporates a contraction function founded on piecewise projection. This contraction function is defined as follows:
[…]
where $||\cdot||_{\infty}$ denotes the infinity norm, defined as $||x||_{\infty}=\text{max}_{j}|x_{j}|$.
4.2. Voxels to mesh conversion
Based on the above high-resolution voxel grid, we employ Marching Cubes (Lorensen and Cline, 1998) to extract an explicit mesh. However, surfaces derived solely from density information in volume rendering often exhibit roughness and are unsuitable for direct use in hybrid representation. Inspired by NeRF2Mesh (Tang et al., 2023), we adopt a synergistic approach involving differentiable rendering and surface operations to refine the explicit mesh.
The objective is to convert a subset of voxels that demonstrate similar rendering quality in both voxel and explicit mesh representations into a refined mesh. Typically, these converted voxels correspond to regions in the scene characterized by uncomplicated geometry and texture. The surface training specifically entails the coarse selection of voxels and subsequent surface refinement using differentiable rendering techniques.
Coarse selection of voxels. By analyzing the rendering results of NeRF2Mesh (Tang et al., 2023) (Figure 5) and BakedSDF (Yariv et al., 2023), we found that using mesh representation to reconstruct the unbounded background region far from the camera presents a formidable challenge. Achieving rendering quality comparable to voxels in this remote area is inherently demanding. Moreover, influenced by the contraction function, triangles extracted from voxels in the contraction region tend to enlarge and may break apart as they extend away from the camera. The presence of large triangles can adversely affect rendering quality, particularly on WebGL platforms, while introducing subdivision for these large triangles to enhance rendering quality incurs additional storage overhead.
Thus, we only convert voxels with surfaces located in the region near the camera into mesh. Criteria such as edge length and the number of triangles in isolated patches guide the decision on which explicit surfaces to retain. By default, we preserve the cube area with $90\%$ of the length of the bounding box in outdoor scenes. This conversion strategy, depicted in Figure 3, contrasts with indiscriminate conversion of voxels containing implicit surfaces into triangular meshes, which results in diminished rendering quality. So these explicit surfaces are further refined in the next differentiable appearance optimization.
Figure 3. The voxels to mesh conversion without (w/o) coarse selection of voxels obtains worse rendering results than with (w/) coarse selection, as shown by the diffuse of Vosh for voxel and mesh components respectively.
Differentiable surface refinement. We perform differentiable rasterization operations on the mesh to obtain intersections between rays and surfaces. Appearance information is obtained by querying the voxel representation based on the corresponding spatial positions. We also employ nvdiffrast (Laine et al., 2020) for differentiable rendering to optimize the appearance information of the mesh, while simultaneously tracking the rendering error corresponding to each face for surface optimization. It is worth noting that we do not enforce the mesh to represent the entire scene. Therefore, during differentiable rendering, we only calculate the loss for pixels corresponding to rays that intersect with the surface. We also perform face merging and mid-point subdivision operations based on the normal change of the faces and rendering errors. In contrast to NeRF2Mesh (Tang et al., 2023), which solely relies on rendering errors for operations, we additionally utilize the local normal change rate of faces as an optimization criterion. This enhances face density in regions with complex geometry and reduces storage overhead for faces in areas with simple geometry.
Following the conversion from voxels to mesh, we have obtained an initial hybrid representation that includes both voxels and mesh. Nevertheless, recognizing significant potential for optimization in rendering quality and volume occupancy for both representations, we are committed to further refining the hybrid representation through further fine-tuning.
Figure 4. The proposed hybrid rendering (c) integrates volume rendering (a) and surface rendering (b) for the Vosh.
4.3. Optimization
After the mesh conversion and optimization stage, we obtain an explicit surface suitable for representing regions in the scene with simple geometry and textures, as well as a optimized high-resolution voxel grid. By combining these two components, we can obtain a coarse hybrid representation with a significant amount of redundancy between the components. This may result in low rendering efficiency and potential rendering conflicts. So we further perform hybrid rendering optimization and voxel adjustment based on the components for a more compact and efficient representation.
Compared to volume rendering, surface rendering only considers the intersection points between rays and surface, significantly reducing the time cost associated with ray marching. This inspires us to confine the ray marching in volume rendering to occur only before the rays hit the surface. Simultaneously, performing volume rendering between the viewpoint and the surface also provides possibilities for enhancing surface rendering quality. Therefore, we straightforwardly combine volume rendering and surface rendering, treating the surface points obtained through rasterization as the final points reached by ray marching in volume rendering. The hybrid rendering can be expressed by the following three equations:
(4) $\mathrm{\mathbf{C}}_{vosh}=\mathrm{\mathbf{C}}_{vosh,d}+\text{MLP}(\mathrm{\mathbf{C}}_{vosh,d},\mathrm{\mathbf{F}_{vosh},\mathrm{\mathbf{d}}}),$
(5) $\mathrm{\mathbf{F}}_{vosh}=\sum_{i}w_{i}\mathrm{\mathbf{f}}+w_{m}\mathrm{\mathbf{f}}_{m},$
(6) $\mathrm{\mathbf{C}}_{vosh,d}=\sum_{i}w_{i}\mathrm{\mathbf{c}}_{d,i}+w_{m}\mathrm{\mathbf{c}}_{d,m},$
where $\mathrm{\mathbf{C}}_{vosh}$, $\mathrm{\mathbf{C}}_{vosh,d}$ and $\mathrm{\mathbf{F}}_{vosh}$ represent the color, diffuse color, and view-dependent features after hybrid rendering, respectively. $w_{m}=(1-\sum_{i}w_{i})$ represents the rendering weight for the mesh, which is the remaining portion of the cumulative voxel weights. $\mathrm{\mathbf{c}}_{d,m}$ and $\mathrm{\mathbf{f}}_{m}$ represent the diffuse color and view-dependent features at the intersection point where the ray hits the surface. Three different rendering methods are shown in Figure 4.
Based on the optimized mesh, we attempt to dynamically adjust the voxel grid, forming a high-quality hybrid representation. Since the trained mesh already possesses the capability to represent regions with simple geometry and textures, we need to consider how to simplify the voxel representation for the entire scene. We focus on retaining voxels in regions with complex geometry and textures to better complement the rendering of the hybrid representation with the mesh. Inspired by PlenOctrees (Yu et al., 2021), we propose a voxel adjustment loss based on surface mesh weights. Specifically, we calculate the rendering weight of mesh $w_{m}$ for rays hitting the surface and force it close to 1. The implementation is as follows:
(7) $\mathcal{L}_{voxel}=\lambda_{voxel}\frac{1}{K}\sum_{k}(1-\text{exp}(w_{m}-1)),$
where $\lambda_{voxel}$ represents the weight of the voxel adjustment loss, and $K$ is the number of samples taken in the rays hitting the surface.
The number of retained voxels is an important factor in determining the balance between rendering quality and speed. Concretely, retaining more voxels can enhance rendering quality but slow down rendering speed, while retaining fewer voxels can accelerate rendering speed but compromise rendering quality. Setting different voxel adjustment weights to limit the volumetric density during ray marching is one of the methods to achieve different trade-offs between speed and quality. Additionally, voxels and mesh may overlap at the same spatial location, potentially affecting rendering quality and the volume of exported assets. To address this issue, we first calculate the occupancy grid of the meshes under the training poses. During voxel adjustment, we ignore the voxels in the mesh-occupied grid. We also adjust the ratio of sparse voxels in the hybrid representation by setting different resolutions $r_{mesh}$ for the mesh-occupancy grid.
Table 1. Quantitative evaluation of the proposed method and SOTA methods with 3 outdoor scenes and 4 indoor scenes from Mip-NeRF 360 dataset (Barron et al., 2022). The methods that can render in real-time on mobile phones are highlighted.
Outdoor Indoor GTX 1660Ti iPhone 14 Pro
Methods Representations PSNR ↑ SSIM ↑ LPIPS ↓ PSNR ↑ SSIM ↑ LPIPS ↓ FPS @ 1920*1080 FPS @ 344*707
NeRF MLP 21.46 0.458 0.515 26.84 0.790 0.370 - -
Mip-NeRF 360 MLP 25.92 0.747 0.244 31.72 0.917 0.179 - -
Instant-NGP Voxel 23.90 0.648 0.369 29.47 0.877 0.273 - -
SNeRG++${}_{1024}$ Voxel 24.11 0.625 0.377 27.32 0.802 0.276 9.0 -
MERF Voxel + Plane 24.41 0.678 0.298 27.80 0.855 0.271 9.7 25.3
BakedSDF Mesh 23.40 0.577 0.351 27.20 0.845 0.300 131.0 -
MobileNeRF Mesh 22.90 0.524 0.463 25.75 0.757 0.453 92.7 153.3
NeRF2Mesh Mesh 22.74 0.497 0.463 25.20 0.721 0.340 231.0 210.0
Vosh-Light Voxel + Mesh 23.63 0.586 0.392 27.02 0.792 0.294 79.7 70.0
Vosh-Base Voxel + Mesh 23.74 0.592 0.389 27.07 0.791 0.291 38.0 33.7
5. Real-Time Rendering
The real-time rendering based on Vosh is realized as a JavaScript 3D web application, leveraging the capabilities of the MERF real-time renderer. The rendering process is also hybrid and structured into two passes: the rasterization pass and the ray marching pass. Both passes are employed to make rendering result consistency between training on the GPU and inference on the web.
In the rasterization pass, meshes are uniformly shaded. The fragment shader takes charge of sampling appearance textures, incorporating diffuse color and view-dependent feature. The output of the rasterization pass includes corresponding diffuse, view-dependent features and depth map. These output is then used in the subsequent ray marching pass.
In the ray marching pass, we also employ a multi-resolution hierarchical structure for the occupancy grid introduced in SNeRG (Hedman et al., 2021) to mark the position of occupied voxels. This structure creates binary mask images in multiple resolution levels through max-pooling. We use the multi-resolution occupancy grid to perform spatial skips during ray marching. Simultaneously, the depth map obtained in the rasterization pass further reduces the sampling distance.
For each spatial coordinate after ray marching, we compare the transformed depth with the depth map from the rasterization pass to determine whether the ray has intersected the surface position. If the ray reaches the surface of the mesh, the marching stops, and the accumulated result of ray marching is weightedly combined with the appearance result from the rasterization pass. Leveraging the surface depth information from the rasterization pass proves instrumental in significantly diminishing the number of required sampling points during ray marching, thereby effectively reducing the computational cost associated with volume rendering.
6. Experiments
We have implemented real-time view synthesis using the proposed hybrid representation Vosh, and carried out comprehensive evaluations encompassing both rendering quality and speed. We also compare our method with state-of-the-art (SOTA) methods, like offline classical methods (NeRF (Mildenhall et al., 2021), Mip-NeRF 360 (Barron et al., 2022), Instant-NGP (Müller et al., 2022)) and real-time methods (Mobile-Nerf (Chen et al., 2023), NeRF2Mesh (Tang et al., 2023), BakedSDF (Yariv et al., 2023), MERF (Reiser et al., 2023), SNeRG++ (Hedman et al., 2021)). As done by SOTA methods, we used the challenging unbounded dataset Mip-NeRF 360 (Barron et al., 2022) for evaluation. Next, we elaborate the details of the experimental results.
Figure 5. The rendering results and zoomed-in images obtained by our method, as well as some SOTA methods based on voxels or mesh representation.
6.1. Rendering quality comparison
We conduct both quantitative and qualitative evaluation of rendering quality for view synthesis obtained by the proposed method and SOTA methods. Quantitative evaluation uses the metrics of PSNR, SSIM (Wang et al., 2004) and LPIPS (Zhang et al., 2018) for the assessment of rendering quality. Table 1 shows the statistics of the quantitative evaluation of different methods.
Here, we classify these methods based on their ability to achieve real-time rendering on mobile devices like laptops and mobile phones. Methods such as SNeRG, MERF, and BakedSDF can only attain real-time rendering on select laptops with robust GPUs. It can be seen there exists an obvious gap between representations based on voxels and meshes. While BakedSDF narrows the quality gap with its millions of triangles, its substantial memory consumption makes it impractical for deployment on mobile phones.
Vosh demonstrates superior rendering quality in real-time rendering on mobile devices, even surpassing some methods that cannot achieve real-time rendering on mobile devices. For the outdoor scenes, Vosh enhances rendering quality over mesh-based methods by leveraging sparse voxels through the use of a contraction function (see Figure 5, 6 and 7). However, for the indoor scenes, methods have limitations in the representational capacity of tiny view-dependent MLPs. While Vosh achieves real-time rendering on mobile devices by employing a tiny view-dependent MLP, it has slightly lower rendering quality in indoor scenes compared to MLP-based methods.
We provide a qualitative comparison with SOTA methods based on voxels (Reiser et al., 2023) and mesh (Tang et al., 2023) on the examples in Figure 5. We select one distant view and one close-up view for local magnification for better exhibition. The mesh-based methods often lack the ability to represent background regions far from the camera location, whereas voxel-based methods necessitate higher computational and storage costs to achieve enhanced rendering quality. Contrarily, our method improves rendering speed while essentially maintaining the rendering quality comparable to that of voxel-based methods. We refer the reader to the companion video for visual demonstrations of the results.
Comparison with VMesh. It is worth noting that VMesh (Guo et al., 2023) is also built on the hybrid representation of volume and mesh, but primarily designed for rendering individual objects or scenes with boundaries, posing challenges in handling unbounded scenes like the examples in Figure 5 and 7. Nonetheless, our method produces better rendering results than VMesh, even for the synthetic bounded object as shown in Figure 6.
Figure 6. The rendering results obtained by ours and VMesh (Guo et al., 2023).
6.2. Rendering speed comparison
We evaluated the performance of the methods applicable for rendering on mobile devices, including SNeRG++ (Hedman et al., 2021), MERF (Reiser et al., 2023), BakedSDF (Yariv et al., 2023), MobileNeRF (Chen et al., 2023), and NeRF2Mesh (Tang et al., 2023). For SNeRG++ and MERF, we follow their configurations with progressive upsampling turned off. We conduct testing on the iPhone 14 Pro equipped with the A16 chip with a resolution of 344*707 and on the Legion Y7000P 2019, which features NVIDIA GeForce GTX 1660Ti (Mobile version, 80W) and operates with a resolution of 1920*1080.
Rendering speed tests for all comparative methods were conducted on the Mip-Nerf 360 dataset (Barron et al., 2022) in three outdoor scenes, using identical device settings, resolutions, and viewpoints. As shown on the right side of Table 1, our method can run with the real-time frame rate (¿30 FPS) on both iPhone 14 Pro and GTX 1660Ti. Particularly on the 1660Ti, our method demonstrates a considerable speed improvement compared to MERF and SNeRG. The latter two methods experience slow rendering speeds due to the necessity to sample each alive voxel traversed by every ray. BakedSDF is difficult to render on mobile phones and laptops with limited VRAM due to its substantial assets. MobileNeRF and NeRF2Mesh achieve rendering speeds that reach the maximum refresh rates of their respective devices (60Hz for Safari on iPhone 14 Pro and 144Hz for Legion Y7000P 2019). We make statistics on their potential rendering speeds using overlapping multiple passes.
6.3. Ablation Study
We conduct ablation experiments on the voxels to mesh conversion and voxel adjustment parts of the hybrid representation in Table 2. The assessment involves a quantitative comparison, considering both rendering quality and speed.
The ‘full bound’ row refers to the results of the variant that do not use coarse selection of voxels in the voxels to mesh conversion. The experimental results show that converting all voxels containing surfaces to mesh not only reduces the rendering quality of the final hybrid representation, but also significantly increases the storage of the assets.
In the voxel adjustment stage, we compare the complete model with variants of Vosh-Base, incorporating different loss weights or resolutions for mesh occupancy grids. The rows with $\lambda_{voxel}$ show the results of different hyper-parameter weights of the loss function for voxel adjustment. The rows with $r_{mesh}$ indicate the results of using mesh occupancy grids with different resolutions. For the Vosh-Base model, we use $\lambda=0.001$ and $r_{mesh}=128$ as hyper-parameters. For the Vosh-Light model, we use $\lambda=0.1$ and $r_{mesh}=32$ as hyper-parameters. Experimental results suggest that increasing the weight of voxel adjustment loss or reducing the resolution of the mesh occupancy grid aids in compressing the number of voxels, leading to speed benefits at the expense of rendering quality.
Table 2. The ablation study in the construction of hybrid representation.
Variants PSNR ↑ SSIM ↑ LPIPS ↓ FPS ↑
Vosh-Base 23.74 0.592 0.389 33.7
full bound 23.49 0.583 0.395 36.0
$\lambda_{voxel}=0$ 23.73 0.591 0.389 30.0
$\lambda_{voxel}=0.01$ 23.72 0.590 0.389 38.7
$\lambda_{voxel}=0.1$ 23.69 0.590 0.390 46.7
$r_{mesh}=0$ 23.80 0.593 0.386 32.0
$r_{mesh}=1024$ 23.79 0.593 0.387 32.0
$r_{mesh}=64$ 23.70 0.590 0.389 37.7
Vosh-Light 23.63 0.586 0.392 70.0
6.4. Limitations
Although the proposed hybrid representation Vosh brings valuable advantages to real-time view synthesis, it is not without limitations. We adopt the same view-dependent MLP as SNeRG and MERF, consequently sharing some similarities with their drawbacks. Notably, it faces challenges such as the incapability to model view-dependent colors for translucent objects, and represent extremely large scenes or objects with complex reflections. These limitations, in turn, may result in potential degradation of mesh optimization quality, thereby impacting the rendering quality based on the hybrid representation.
7. Conclusion
We have presented a hybrid representation Vosh, composed of both voxels and mesh, for NeRF-based real-time view synthesis. By capitalizing on the strengths of voxel and mesh components in our representation, we attain a controllable balance between rendering quality and speed. Experimental evidence demonstrates that our approach outperforms state-of-the-art methods, showcasing exceptional flexibility for deployment across a range of mobile devices.
In our future work, we intend to explore more efficient data structures for organizing and storing the hybrid representation, aiming to optimize memory consumption and further enhance rendering efficiency. Additionally, we are considering more compact integration of classic representations to achieve fast and high-quality novel view synthesis specifically tailored for mobile devices.

### v2

@@
-Vosh: Voxel-Mesh Hybrid Representation for Real-Time View Synthesis
-Abstract.
-The neural radiance field (NeRF) has emerged as a prominent methodology for synthesizing realistic images of novel views.
+Voxel-Mesh Hybrid Representation For Real-Time View Synthesis By Meshing Density Field
+Abstract
+The neural radiance fields (NeRF) have emerged as a prominent methodology for synthesizing realistic images of novel views.
 While neural radiance representations based on voxels or mesh individually offer distinct advantages, excelling in either rendering quality or speed, each has limitations in the other aspect.
-In response, we propose a pioneering hybrid representation named Vosh, seamlessly combining both voxel and mesh components in hybrid rendering for view synthesis.
-Vosh is meticulously crafted by optimizing the voxel grid of NeRF, strategically with selected voxels replaced by mesh.
+In response, we propose a hybrid representation named Vosh, seamlessly combining both voxel and mesh components in hybrid rendering for view synthesis.
+Vosh is meticulously crafted by optimizing the voxel grid based on neural rendering, strategically meshing a portion of the volumetric density field to surface.
 Therefore, it excels in fast rendering scenes with simple geometry and textures through its mesh component, while simultaneously enabling high-quality rendering in intricate regions by leveraging voxel component.
 The flexibility of Vosh is showcased through the ability to adjust hybrid ratios, providing users the ability to control the balance between rendering quality and speed based on flexible usage.
-Experimental results demonstrates that our method achieves commendable trade-off between rendering quality and speed, and notably has real-time performance on mobile devices.
-Neural Radiance Field, View Synthesis, Real-Time Rendering.
-journal: TOGcopyright: noneccs: Computing methodologies Reconstructionccs: Computing methodologies Neural networksccs: Computing methodologies Volumetric models
-Figure 1.
-Left: The hybrid representation Vosh, combines both voxel component (highlighted with pseudo colors) and mesh component (indicated by gray triangles) in hybrid rendering.
+Experimental results demonstrate that our method achieves commendable trade-off between rendering quality and speed, and notably has real-time performance on mobile devices.
+The interactive web demo and code are available at https://zyyzyy06.github.io/Vosh.
+Index Terms:
+Neural Radiance Fields, View Synthesis, Real-Time Rendering.
+Figure 1: Left: The hybrid representation Vosh, combines both voxel component (highlighted with pseudo colors) and mesh component (indicated by gray triangles) in hybrid rendering.
 Right: The proposed method facilitates real-time view synthesis on various mobile devices, e.g., laptops and mobile phones.
-1. Introduction
-The novel view synthesis based on the Neural Radiance Field (NeRF) leverages neural networks to model three-dimensional scenes and generates images from specified viewpoints (Mildenhall et al., 2021).
+I Introduction
+The novel view synthesis based on the Neural Radiance Fields (NeRF) leverages neural networks to model three-dimensional scenes and generates images from specified viewpoints [1].
 In general, NeRF-based approaches demonstrate proficiency in producing high-quality rendering for novel views with sparse inputs, effectively handling materials with specular and translucent properties.
-This versatility positions them for diverse applications in autonomous navigation (Kwon et al., 2023), virtual reality/augmented reality (Li et al., 2022a), and beyond (Sun et al., 2022b; Jiang et al., 2023).
+This versatility positions them for diverse applications in autonomous navigation [2], virtual reality/augmented reality [3], and beyond [4, 5].
 Despite these merits, current methods face challenges in achieving a balance between high-quality rendering and real-time rendering, particularly when deployed on mobile devices.
-The rendering speed of neural radiance fields is significantly influenced by the chosen representation, with notable implications for sampling requirements and query costs (Yu et al., 2021).
+The rendering speed of neural radiance fields is significantly influenced by the chosen representation, with notable implications for sampling requirements and query costs [6].
 Typically, the original NeRF uses deep multilayer perceptrons (MLPs) to model neural radiance fields, relying on dense sampling for querying to approximate integrals in volume rendering.
 Unfortunately, this approach typically causes excessively long rendering times.
-A substantial amount of effort has explored alternatives such as implicit voxel representations (Sun et al., 2022a; Hedman et al., 2021; Liu et al., 2020; Fridovich-Keil et al., 2022; Müller et al., 2022; Reiser et al., 2021) and explicit mesh representations (Chen et al., 2023; Tang et al., 2023; Rakotosaona et al., 2023; Yariv et al., 2023), aiming to exclude irrelevant regions for computation.
-These methods reduce the number of sampling points and employ smaller MLPs to expedite rendering.
-However, even though approaches based on implicit voxels can achieve high-quality rendering in complex scenes, the dense sampling in non-empty regions is prohibitively expensive.
-Methods based on mesh representation have demonstrated significantly accelerated rendering speeds, enabling real-time rendering on common mobile devices.
-However, the challenge lies in the reconstruction of scene-level mesh structure, particularly in areas with intricate geometry and textures (see Figure 1), such as thin structures or undetermined regions like backgrounds (Reiser et al., 2023).
+A substantial amount of effort has explored alternatives such as voxel representations [7, 8, 9, 10, 11, 12] and mesh representations [13, 14, 15, 16], aiming to exclude irrelevant regions for computation.
+These methods reduce the number of sampling points and employ smaller MLPs or no MLPs to expedite rendering.
+However, even though approaches based on voxels can achieve high-quality rendering in complex scenes, the dense sampling in non-empty regions is prohibitively expensive.
+3DGS [17], with its Gaussian representation and Splatting rendering method, also reduces the number of sampled points in rays, achieving real-time rendering of high-resolution scenes.
+But its high memory usage and platform-related sorting strategies hinder its widespread applications, especially on mobile devices.
+The methods based on mesh representation have demonstrated significantly accelerated rendering speeds, enabling real-time rendering on common mobile devices.
+However, the challenge lies in the reconstruction of scene-level mesh structure, particularly in areas with intricate geometry and textures (see Figure 1), such as thin structures or undetermined regions like backgrounds [18].
 It is observed that meshes can match the rendering quality akin to voxels in regions with simple geometry and textures, as well as offering highly efficient rendering.
 Therefore, a promising solution involves employing hybrid representations for reconstruction towards both high-quality rendering and fast speed.
-Existing works on hybrid representations (Wang et al., 2023; Guo et al., 2023), focus primarily on object or proximal region reconstruction, thus limiting flexibility to represent complex scenes and control over different parts of the hybrid representation.
-Alternatively, we leverage the strengths of voxels and mesh, to design a hybrid representation towards a better balance between quality and speed, especially enabling high-quality rendering on mobile devices (see Figure 1).
+Existing works on hybrid representations [19, 20], focus primarily on object or proximal region reconstruction, thus abandoning the reconstruction ability of volume representation for complex scenes, such as background regions lacking observation in unbounded outdoor environments.
+VMesh [20] mainly relies on the signed distance function for the scene reconstruction.
+However, it may result in low-quality reconstruction and is space-consuming for complex background areas of unbounded scenes.
+We utilize implicit volumetric density field, which can recover more comprehensive information and be applied in both bounded and unbounded scenes.
+By partially meshing the density field, a more expressive mesh representation can be obtained.
+Then we leverage the strengths of voxels and mesh, to design a hybrid representation towards a better balance between quality and speed, especially enabling high-quality rendering on mobile devices (see Figure 1).
 In summary, the main contribution of this paper is a new hybrid representation namely Vosh (VOxel-meSH), by combing voxels and mesh for NeRF-based real-time view synthesis.
-It also allows for a dynamic control of the balance between voxel and mesh components by voxel adjustment.
+It also allows for a dynamic control of the balance between voxel and mesh components.
 Such flexibility results in a versatile spectrum of NeRF representations, catering to the varied computational capabilities that facilitate real-time view synthesis across different devices.
-Experimental results exhibit the super performance of our method compared to state-of-the-art methods.
-Figure 2.
-An overview of the proposed methodology.
+Experimental results exhibit the superior performance of our method compared to state-of-the-art methods.
+II Related Work
+There are numerous published approaches to view synthesis and NeRF, and we refer interested readers to [21, 22, 23] for a comprehensive overview.
+In this context, we focus on the real-time view synthesis and NeRF representations most related to our method.
+II-A Real-time view synthesis
+NeRF [1] as a well-known technology for view synthesis, has been the focus of numerous endeavors aimed at accelerating its rendering speed while maintaining high quality of rendered images.
+SNeRG [8] achieves fast rendering without CUDA, using sparse voxel representations and a compact neural network for deferred rendering.
+Plenoxels [10] and PlenOctrees [6] expedite ray marching with sparse voxel representation, refining voxels near object surfaces.
+MERF [18] enhances rendering speed and spatial occupancy by replacing pure voxel grids with a low-resolution voxel grid and a high-resolution triplane projection representation [24].
+These methods mostly excel on high-end GPUs, whereas real-time rendering on mobile devices still remains a challenge.
+MobileNeRF [13] achieves fast rendering by constraining NeRF to triangles and using rasterization.
+NeRF2Mesh [14] uses NeRF to initialize geometric structures, iteratively refining mesh attributes to improve rendering quality.
+BakedSDF [16] optimizes a blended neural signed distance function representation based on VolSDF [25], enabling real-time rendering of complex scenes on a high-quality mesh.
+In order to represent complex backgrounds, BakedSDF uses a contracted space containing Eikonal loss which might have negative impact on the quality of the foreground as claimed in [26].
+While mesh-based methods offer fast rendering, they suffer quality loss compared to volume-based methods, especially for unbounded scenes with complex backgrounds.
+3D Gaussian Splatting [17] creates a 3D Gaussian representation around each point starting from a point cloud model, utilizing the Splatting method for real-time rendering of high-resolution scenes.
+To achieve real-time training and rendering, 3DGS relies on CUDA-specific APIs.
+So original 3DGS is difficult to render on most mobile devices.
+While recent re-implementations seem to render 3DGS models on commodity hardware, they rely on approximations to sort order and view-dependency whose impact on quality and speed has not yet been evaluated [27].
+In contrast, our hybrid representation combines the strengths of voxel and mesh representations, allowing high-quality rendering of complex scenes on mobile devices.
+Recently, 3DGS-based compression methods have emerged to effectively reduce the number of Gaussians.
+Eagles [28] proposes to use quantized embeddings and a coarse-to-fine training strategy to significantly reduce memory storage overhead.
+F-3DGS [29] represents dense clusters of Gaussians with significantly fewer Gaussians through efficient factorization.
+While these methods have successfully enhanced 3DGS in terms of storage efficiency and rendering speed, they still require as much or even more VRAM in the rendering compared to the original 3DGS as reported in these papers, which are still not suitable for the use on mobile devices.
+II-B Hybrid scene representation
+Combining explicit mesh representation with implicit fields paves a new way to NeRF-based processing, such as NeRF editing and rendering acceleration.
+NeuMesh [30] encodes a neural implicit field with disentangled geometry and texture on mesh vertices, enabling a range of neural rendering-based editing operations.
+EyeNeRF [31] integrates an explicit surface with an implicit volume representation to model the human eye.
+Mesh-Aware-RF [32] embeds a known polygonal mesh into NeRF, employing ray tracing and ray marching alternately to update radiance along rays.
+The hybrid representation like Adaptive Shells [19] and VMesh [20] has been used in view synthesis.
+Adaptive Shells confine volume rendering to a narrow band around the object for effective NeRF rendering.
+VMesh, closely related to our work, uses a textured mesh and auxiliary sparse voxels but differs by employing the signed distance function (SDF) and volume density field to simultaneously learn the bounded scenes or objects.
+Both methods use NeuS [33] as the implementation for SDF.
+NeuS learns neural implicit surfaces through volume rendering.
+For the unbounded scenes with backgrounds, NeuS needs to use NeRF for background modeling that is difficult to represent by itself, which is also a hybrid representation.
+It should be noted that VMesh most closely resembling our method in this paper, relies on SDF with density field to capture scene information, while our method only uses the density field.
+Specifically, when using SDF to reconstruct complex areas like cluttered backgrounds in unbounded scenes, it might lead to low-quality reconstruction and significant space occupation [16].
+Due to this limitation, VMesh is primarily suited for single objects or bounded scenes, whereas our density-field-based method effectively manages a wide range of scenarios, including unbounded outdoor scenes.
+Besides, VMesh requires manual tuning of multiple loss weights for each scene to achieve optimal surface-volume separation as indicated in the paper [20].
+Without these adjustments, only a single representation may be optimized.
+In contrast, our method based on the density field does not rely on scene-specific hyper-parameters, enabling the high-quality hybrid representation for various scenes.
+Figure 2: An overview of the proposed methodology.
 The training phase starts from grid training for obtaining initial voxels.
-Then, a portion of voxels becomes mesh in the voxels to mesh conversion.
-Subsequently, the combination of voxels and mesh are optimized through hybrid rendering and voxels adjustment to obtain the final hybrid representation Vosh.
+Then, a portion of initial volumetric density field are meshed.
+Subsequently, the combination of voxels and mesh are optimized through hybrid rendering and voxel pruning to obtain the final hybrid representation Vosh.
 The inference phase realizes real-time hybrid rendering with Vosh even on mobile phones.
-2. Related Work
-There are numerous published approaches to view synthesis and NeRF, and we refer interested readers to (Tewari et al., 2022; Gao et al., 2022) for a comprehensive overview.
-In this context, we focus on the real-time view synthesis and NeRF representations most related to our method.
-2.1. Real-time view synthesis
-NeRF (Mildenhall et al., 2021) as a well-known technology for view synthesis, has been the focus of numerous endeavors aimed at enhancing its rendering speed while maintaining high-quality rendering.
-SNeRG (Hedman et al., 2021) achieves fast rendering without CUDA, using sparse voxel representations and a compact neural network for deferred rendering.
-Plenoxels (Fridovich-Keil et al., 2022) and PlenOctrees (Yu et al., 2021) expedite ray marching with sparse voxel representation, refining voxels near object surfaces.
-MERF (Reiser et al., 2023) enhances rendering speed and spatial occupancy by replacing pure voxel grids with a low-resolution voxel grid and a high-resolution triplane projection representation (Chan et al., 2022).
-These methods mostly excel on high-end GPUs, whereas real-time rendering on mobile devices still remains a challenge.
-MobileNeRF (Chen et al., 2023) achieves fast rendering by constraining NeRF to triangles and using rasterization.
-NeRF2Mesh (Tang et al., 2023) uses NeRF to initialize geometric structures, iteratively refining mesh attributes to improve rendering quality.
-BakedSDF (Yariv et al., 2023) optimizes a blended neural signed distance function representation, enabling real-time rendering of complex scenes on a high-quality mesh.
-While mesh-based methods offer fast rendering, they may suffer quality loss compared to voxel-based methods.
-3D Gaussian Splatting (Kerbl et al., 2023) creates a 3D Gaussian representation around each point starting from a point cloud model, utilizing the Splatting method for real-time rendering of high-resolution scenes.
-But it suffers elongated and popping artifacts, and high memory usage for deployment on mobile phones.
-In contrast, our hybrid representation combines the strengths of voxel and mesh representations, allowing high-quality rendering of complex scenes on mobile devices.
-2.2. Hybrid Scene Representation
-Combining explicit mesh representation with implicit fields introduces functionalities such as NeRF editing and rendering acceleration.
-NeuMesh (Yang et al., 2022) encodes a neural implicit field with disentangled geometry and texture on mesh vertices, enabling a range of neural rendering-based editing operations.
-EyeNeRF (Li et al., 2022b) integrates an explicit surface with an implicit volume representation to model the human eye.
-Mesh-Aware-RF (Qiao et al., 2023) embeds a known polygonal mesh into NeRF, employing ray tracing and ray marching alternately to update radiance along rays.
-The hybrid representation like Adaptive Shells (Wang et al., 2023) and VMesh (Guo et al., 2023) has been used in view synthesis.
-Adaptive Shells confines volume rendering to a narrow band around the object for effective NeRF rendering.
-VMesh, closely related to our work, uses a textured mesh and auxiliary sparse voxels but differs by employing SDF field and volume density field to simultaneously learn the entire scene.
-However, VMesh relies on gradient descent optimization that lacks controllability and flexibility.
-Indeed, VMesh mainly focuses on rendering individual objects or bounded scenes, while the proposed method in this paper can handle more scenes, such as unbounded outdoor scenes.
-3. Method Overview
+III Method Overview
 We start with an overview of our methodology to establish a context for the core algorithms described in the next section.
@@
 In the inference phase, our method achieves real-time rendering based on the baked hybrid representation assets.
-Concretely, the training phase consists of three stages: grid training, voxels to mesh conversion and optimization.
+Concretely, the training phase consists of three stages: grid training, meshing density field and optimization.
 In the grid training stage, we train a voxel grid model through volume rendering.
-Then, leveraging the trained voxel grid, we extract an explicit surface and train the surface using differentiable surface rendering for both appearance and geometry optimization.
+Then, we determine the region of meshing in the density field based on the difference in rendering quality between the voxel grid and the potential surfaces it contains.
 This converts the initial voxel grid into a combination of voxel and mesh components.
-Finally, the voxels and mesh are further optimized through a training process involving hybrid rendering and voxels adjustment to obtain the desired hybrid representation.
+Finally, the voxels and mesh are further optimized through a training process involving hybrid rendering and voxel pruning to obtain the desired hybrid representation.
 Once training is complete, we export assets containing sparse voxels and textured meshes.
-Subsequently, real-time view synthesis on mobile devices is implemented using WebGL.
+Subsequently, in the inference phase, we implement the rendering process of Vosh based on WebGL, enabling real-time view synthesis on mobile devices.
 Next, we elaborate the algorithm details in the training and inference phases.
-4. Hybrid Representation construction
+IV Hybrid Representation construction
 The construction of a hybrid representation in the training phase commences with the initialization of a volume comprising solely of voxels.
-This volume is then subject to a conversion process wherein a portion of the voxels is replaced with an explicit mesh according to rendering quality on par with the original voxel representation.
-4.1. Grid training for initial voxels
+Then, by calculating the difference in rendering quality between the volume and its underlying surface, the density field corresponding to regions with similar rendering quality are meshed to replace the volumetric density values corresponding to the same positions.
+IV-A Grid training for initial voxels
 Initially, the entire scene is reconstructed using a high-resolution voxel grid.
 It retains sufficient geometric and appearance information essential for subsequent mesh extraction and hybrid representation.
-Here, we employ SNeRG++ (an improved version of SNeRG (Hedman et al., 2021) in MERF (Reiser et al., 2023)) as the benchmark for our voxel grid, due to its capability to effectively model unbounded scenes and generate real-time rendering assets.
-Concretely, SNeRG++ is built upon the volume rendering principles of vanilla NeRF, which involves sampling along the direction of a ray originating from the ray origin $o$ along the direction $d$ to obtain sample points $\mathrm{\mathbf{x}}_{i}=o+t_{i}\mathrm{\mathbf{d}}$.
+Here, we employ SNeRG++ (an improved version of SNeRG [8] in MERF [18]) as the benchmark for our voxel grid, due to its capability to effectively model unbounded scenes and generate real-time rendering assets.
+Concretely, SNeRG++ is built upon the volume rendering principles of NeRF, which involves sampling along the direction of a ray originating from the ray origin $\mathbf{o}$ along the direction $\mathbf{d}$ to obtain sample points $\mathrm{\mathbf{x}}_{i}=\textbf{o}+t_{i}\mathrm{\mathbf{d}}$.
 Based on the spatial position of these sample points, it uses the neural radiance field to calculate volumetric density $\sigma_{i}$, resulting in rendering weights $w_{i}$:
-(1) $w_{i}=\alpha_{i}T_{i},\;\;\;\;\;\;\;\;T_{i}=\prod^{i-1}_{j=1}(1-\alpha_{j}),\;\;\;\;\;\;\;\;\alpha_{i}=1-e^{-\sigma_{i}\delta_{i}},$
+$w_{i}=\alpha_{i}T_{i},\;\;\;\;\;\;\;\;T_{i}=\prod^{i-1}_{j=1}(1-\alpha_{j}),\;\;\;\;\;\;\;\;\alpha_{i}=1-e^{-\sigma_{i}\delta_{i}},$ (1)
 where $T_{i}$, $\alpha_{i}$ and $\delta_{i}$ represent the transmittance, opacity and spacing of sampling point $i$, respectively.
-SNeRG++ adopts the concept of deferred rendering, decomposing the appearance information of the neural radiance field into diffuse color $\mathrm{\mathbf{c}}_{d,i}$ and view-dependent features $f_{i}$.
-By cumulatively sampling the appearance information along the ray, as shown in Figure 4a, we obtain the diffuse color $C_{d}$ and view-dependent features $F$ of the ray as
-(2) $\mathrm{\mathbf{C}}_{d}=\sum_{i}w_{i}\mathrm{\mathbf{c}}_{d,i},\;\;\;\;\;\;\;\;\mathrm{\mathbf{F}}=\sum_{i}w_{i}\mathrm{\mathbf{f}}_{i}.$
-Finally, the view-dependent color is calculated by inputting the diffuse color $\mathrm{\mathbf{C}}_{d}$, view-dependent features $F$, and the ray direction $d$ into a miniature MLP.
-The result is then added to the diffuse color $\mathrm{\mathbf{C}}_{d}$ to obtain the final color $C$ of the ray.
-Besides, in order to enhance the modeling of regions situated far from the center within unbounded scenes and facilitate efficient ray marching calculations, SNeRG++ incorporates a contraction function founded on piecewise projection.
-This contraction function is defined as follows:
-[…]
-where $||\cdot||_{\infty}$ denotes the infinity norm, defined as $||x||_{\infty}=\text{max}_{j}|x_{j}|$.
-4.2. Voxels to mesh conversion
-Based on the above high-resolution voxel grid, we employ Marching Cubes (Lorensen and Cline, 1998) to extract an explicit mesh.
-However, surfaces derived solely from density information in volume rendering often exhibit roughness and are unsuitable for direct use in hybrid representation.
-Inspired by NeRF2Mesh (Tang et al., 2023), we adopt a synergistic approach involving differentiable rendering and surface operations to refine the explicit mesh.
-The objective is to convert a subset of voxels that demonstrate similar rendering quality in both voxel and explicit mesh representations into a refined mesh.
-Typically, these converted voxels correspond to regions in the scene characterized by uncomplicated geometry and texture.
-The surface training specifically entails the coarse selection of voxels and subsequent surface refinement using differentiable rendering techniques.
-Coarse selection of voxels.
-By analyzing the rendering results of NeRF2Mesh (Tang et al., 2023) (Figure 5) and BakedSDF (Yariv et al., 2023), we found that using mesh representation to reconstruct the unbounded background region far from the camera presents a formidable challenge.
-Achieving rendering quality comparable to voxels in this remote area is inherently demanding.
-Moreover, influenced by the contraction function, triangles extracted from voxels in the contraction region tend to enlarge and may break apart as they extend away from the camera.
-The presence of large triangles can adversely affect rendering quality, particularly on WebGL platforms, while introducing subdivision for these large triangles to enhance rendering quality incurs additional storage overhead.
-Thus, we only convert voxels with surfaces located in the region near the camera into mesh.
-Criteria such as edge length and the number of triangles in isolated patches guide the decision on which explicit surfaces to retain.
-By default, we preserve the cube area with $90\%$ of the length of the bounding box in outdoor scenes.
-This conversion strategy, depicted in Figure 3, contrasts with indiscriminate conversion of voxels containing implicit surfaces into triangular meshes, which results in diminished rendering quality.
-So these explicit surfaces are further refined in the next differentiable appearance optimization.
-Figure 3.
-The voxels to mesh conversion without (w/o) coarse selection of voxels obtains worse rendering results than with (w/) coarse selection, as shown by the diffuse of Vosh for voxel and mesh components respectively.
+SNeRG++ adopts the concept of deferred rendering, decomposing the appearance information of the neural radiance field into diffuse color $\mathrm{\mathbf{c}}_{d,i}$ and view-dependent features $\mathbf{f}_{i}$.
+By cumulatively sampling the appearance information along the ray, we can obtain the diffuse color $\mathbf{C}_{d}$ and view-dependent features $\mathbf{F}$ of the ray as
+$\mathrm{\mathbf{C}}_{d}=\sum_{i}w_{i}\mathrm{\mathbf{c}}_{d,i},\;\;\;\;\;\;\;\;\mathrm{\mathbf{F}}=\sum_{i}w_{i}\mathrm{\mathbf{f}}_{i}.$ (2)
+Finally, the view-dependent color is calculated by inputting the diffuse color $\mathrm{\mathbf{C}}_{d}$, view-dependent features $\mathbf{F}$, and the ray direction $d$ into a miniature MLP.
+The result is then added to the diffuse color $\mathrm{\mathbf{C}}_{d}$ to obtain the final color $\mathbf{C}$ of the ray.
+Besides, in order to enhance the modeling of regions situated far from the center within unbounded scenes and facilitate efficient ray marching calculations, SNeRG++ incorporates a contraction function defined on piecewise projection.
+IV-B Meshing density field
+The motivation for meshing density field is to convert as many voxels as possible to mesh with minimal quality loss.
+Usually, there is a significant quality gap between voxel and mesh representation, but the distribution corresponding to the gap is unknown.
+We identify spatial regions where the volume and its implicit surface offer comparable rendering quality by measuring the quality difference between the surface after differentiable refinement and initial voxels, as well as meshing the density field within these regions.
+In this way, we can achieve speed gain with minimal quality loss.
 Differentiable surface refinement.
-We perform differentiable rasterization operations on the mesh to obtain intersections between rays and surfaces.
+Based on the above high-resolution voxel grid, we employ Marching Cubes [34] for the density field of initial voxel grid to extract an explicit mesh for evaluating the quality gap between voxel and mesh representation.
+However, surfaces meshed directly from volumetric density field often exhibit roughness that have to be improved for enhancing the rendering quality.
+Inspired by NeRF2Mesh [14], we adopt an ad-hoc scheme involving differentiable rendering and surface operations to refine the explicit mesh firstly.
+Concretely, we perform differentiable rasterization operations on the mesh to obtain intersections between rays and surfaces.
 Appearance information is obtained by querying the voxel representation based on the corresponding spatial positions.
-We also employ nvdiffrast (Laine et al., 2020) for differentiable rendering to optimize the appearance information of the mesh, while simultaneously tracking the rendering error corresponding to each face for surface optimization.
-It is worth noting that we do not enforce the mesh to represent the entire scene.
-Therefore, during differentiable rendering, we only calculate the loss for pixels corresponding to rays that intersect with the surface.
-We also perform face merging and mid-point subdivision operations based on the normal change of the faces and rendering errors.
-In contrast to NeRF2Mesh (Tang et al., 2023), which solely relies on rendering errors for operations, we additionally utilize the local normal change rate of faces as an optimization criterion.
+Similar to NeRF2Mesh [14], we employ nvdiffrast [35] for differentiable rendering to optimize the appearance and the trainable offset of vertices, while simultaneously tracking the rendering error corresponding to each face for surface optimization.
+In addition to rendering errors, we also calculate the normal angle between the adjacent faces of each face for performing mesh refinement and decimation, thereby improving the representation ability of the mesh and reducing memory usage.
+Specifically, for coarse mesh extracted based on initial voxel grid, we use differentiable rendering to learn offsets for each vertex of the mesh.
+Meanwhile, we calculate the rendering error $E_{render}$ and the normal angle between the adjacent faces $A_{normal}$ for each face.
+We set the rendering error threshold $e_{sub}$ and angle threshold $a_{sub}$ for face subdivision based on percentiles:
+$e_{sub}=percentile(E_{render},90),$ (3)
+$a_{sub}=percentile(A_{normal},95).$ (4)
+Similarly, we set the rendering error threshold $e_{dec}$ and angle change threshold $a_{dec}$ for face simplification:
+$e_{dec}=percentile(E_{render},50),$ (5)
+$a_{dec}=percentile(A_{normal},10).$ (6)
+For faces with $E_{render}>e_{sub}$ or $A_{normal}>a_{sub}$, we perform mid-point subdivision [36] to increase face density and enhance representation ability.
+For faces with $E_{render}<e_{dec}$ or $A_{normal}<a_{dec}$, we perform faces simplification [37] to improve rendering efficiency.
+In contrast to NeRF2Mesh [14], which solely relies on rendering errors for operations, we additionally utilize the angle of local normal of faces as an optimization criterion.
 This enhances face density in regions with complex geometry and reduces storage overhead for faces in areas with simple geometry.
-Following the conversion from voxels to mesh, we have obtained an initial hybrid representation that includes both voxels and mesh.
-Nevertheless, recognizing significant potential for optimization in rendering quality and volume occupancy for both representations, we are committed to further refining the hybrid representation through further fine-tuning.
-Figure 4. The proposed hybrid rendering (c) integrates volume rendering (a) and surface rendering (b) for the Vosh.
-4.3. Optimization
-After the mesh conversion and optimization stage, we obtain an explicit surface suitable for representing regions in the scene with simple geometry and textures, as well as a optimized high-resolution voxel grid.
+Figure 3: Visualization of conversion loss $E_{convert}$ based on rendering quality.
+(a) and (b) are the rendering images based on initial voxels and mesh with differentiable surface refinement, respectively.
+(c) is the gray-scale image of conversion loss $E_{convert}$ calculated based on the rendering quality of both representations.
+(d) is to filter out the top $10\%$ $E_{convert}$ in the space.
+Meshing based on rendering quality.
+We can obtain the conversion error field $E_{convert}$ of different representations in the space based on differentiable rendering.
+Specifically, for voxel representation, we allocate rendering errors to corresponding positions based on the rendering weights of different sampling points.
+For mesh representation, the rendering errors at the corresponding position is recorded at the intersection of the surface and rays.
+We use grids to accumulate rendering errors for different representations and calculate the average rendering errors $E_{voxel}$ for voxel representation and $E_{mesh}$ for mesh representation.
+We calculate the L1 distance $E_{convert}=E_{mesh}-E_{voxel}$ of rendering errors in regions where neither is zero as the conversion errors, as shown in Figure 3.
+It is observed in Figure 3 that the main distribution of $E_{convert}$ is in the background area far from the center of cameras.
+This is probably because 1) the reconstruction quality of the central area which is fully observed by cameras is higher enough in mesh representation; 2) contraction function leads to significant loss of information in background voxels.
+So the meshing operation is focused on the background area that lacks sufficient observation and is affected by the contraction function.
+For the scenes based on contraction function in $[-2,2]$, we define the central region as $[-0.5,0.5]$ and the outside as the background region.
+Then, we count voxels with background area and sort them, and also set the threshold in percentiles as $e_{remove}=percentile(E_{convert},p_{remove})$.
+When $p_{remove}$ is set to 0.9, the voxels with their $E_{convert}$ values greater than the threshold $e_{remove}$ is shown in Figure 3(d).
+Finally, we remove the faces converted from voxels located in the background area with $E_{convert}>e_{remove}$.
+The remaining mesh is the result of meshing the regions with small gap in rendering quality in the density field.
+Following the rendering quality based meshing process in density field, we have obtained an initial hybrid representation that includes both voxels and mesh.
+Nevertheless, considering significant potential for optimization in rendering quality and volume occupancy for both representations, we are committed to further refining the hybrid representation through the next optimization.
+Figure 4: The proposed hybrid rendering (c) integrates volume rendering (a) and surface rendering (b) for the Vosh.
+TABLE I: Quantitative evaluation of the proposed method and SOTA methods on 3 outdoor scenes and 4 indoor scenes from Mip-NeRF 360 dataset [38].
+The methods based on hybrid representation are listed separately below.
+means the method is not supported on the tested device.
+Outdoor Indoor Intel Iris Xe 80EU iPhone 14 Pro
+Methods PSNR ↑ SSIM ↑ LPIPS ↓ PSNR ↑ SSIM ↑ LPIPS ↓ FPS @ 800$\times$600 FPS @ 347$\times$600
+NeRF 21.46 0.458 0.515 26.84 0.790 0.370 - -
+Mip-NeRF 360 25.92 0.747 0.244 31.72 0.917 0.179 - -
+Instant-NGP 23.90 0.648 0.369 29.47 0.877 0.273 - -
+Zip-NeRF 26.87 0.807 0.176 32.03 0.925 0.167 - -
+3DGS 26.40 0.805 0.173 30.41 0.920 0.190 - -
+SNeRG++ 24.48 0.646 0.358 27.90 0.837 0.238 22.0 -
+MERF 24.41 0.678 0.298 27.80 0.855 0.271 17.3 25.3
+BakedSDF 23.40 0.577 0.351 27.20 0.845 0.300 21.0 -
+MobileNeRF 22.90 0.524 0.463 25.75 0.757 0.453 59.3 153.3
+NeRF2Mesh 22.74 0.497 0.463 25.20 0.721 0.340 138.7 210.0
+Adaptive Shells 23.17 0.606 0.389 29.19 0.872 0.285 - -
+VMesh - - - - - - - -
+Vosh-Light 23.73 0.601 0.378 27.20 0.802 0.275 78.0 78.7
+Vosh-Base 24.18 0.623 0.364 27.61 0.818 0.261 38.0 34.0
+IV-C Optimization
+After meshing a portion of density field based on rending quality, we obtain an explicit surface suitable for representing regions in the scene with simple geometry and textures, as well as a optimized high-resolution voxel grid.
 By combining these two components, we can obtain a coarse hybrid representation with a significant amount of redundancy between the components.
-This may result in low rendering efficiency and potential rendering conflicts.
-So we further perform hybrid rendering optimization and voxel adjustment based on the components for a more compact and efficient representation.
+This may result in low rendering efficiency and potential rendering artifacts.
+So we further perform hybrid rendering optimization and voxel pruning based on the components for a more compact and efficient representation.
 Compared to volume rendering, surface rendering only considers the intersection points between rays and surface, significantly reducing the time cost associated with ray marching.
@@
 The hybrid rendering can be expressed by the following three equations:
-(4) $\mathrm{\mathbf{C}}_{vosh}=\mathrm{\mathbf{C}}_{vosh,d}+\text{MLP}(\mathrm{\mathbf{C}}_{vosh,d},\mathrm{\mathbf{F}_{vosh},\mathrm{\mathbf{d}}}),$
-(5) $\mathrm{\mathbf{F}}_{vosh}=\sum_{i}w_{i}\mathrm{\mathbf{f}}+w_{m}\mathrm{\mathbf{f}}_{m},$
-(6) $\mathrm{\mathbf{C}}_{vosh,d}=\sum_{i}w_{i}\mathrm{\mathbf{c}}_{d,i}+w_{m}\mathrm{\mathbf{c}}_{d,m},$
+$\mathrm{\mathbf{C}}_{vosh}=\mathrm{\mathbf{C}}_{vosh,d}+\text{MLP}(\mathrm{\mathbf{C}}_{vosh,d},\mathrm{\mathbf{F}_{vosh},\mathrm{\mathbf{d}}}),$ (7)
+$\mathrm{\mathbf{F}}_{vosh}=\sum_{i}w_{i}\mathrm{\mathbf{f}}+w_{m}\mathrm{\mathbf{f}}_{m},$ (8)
+$\mathrm{\mathbf{C}}_{vosh,d}=\sum_{i}w_{i}\mathrm{\mathbf{c}}_{d,i}+w_{m}\mathrm{\mathbf{c}}_{d,m},$ (9)
 where $\mathrm{\mathbf{C}}_{vosh}$, $\mathrm{\mathbf{C}}_{vosh,d}$ and $\mathrm{\mathbf{F}}_{vosh}$ represent the color, diffuse color, and view-dependent features after hybrid rendering, respectively.
@@
 $\mathrm{\mathbf{c}}_{d,m}$ and $\mathrm{\mathbf{f}}_{m}$ represent the diffuse color and view-dependent features at the intersection point where the ray hits the surface.
-Three different rendering methods are shown in Figure 4.
-Based on the optimized mesh, we attempt to dynamically adjust the voxel grid, forming a high-quality hybrid representation.
+Three different rendering procedures are shown in Figure 4.
+Based on the optimized mesh, we attempt to dynamically prune the voxel grid, forming a high-quality hybrid representation.
 Since the trained mesh already possesses the capability to represent regions with simple geometry and textures, we need to consider how to simplify the voxel representation for the entire scene.
 We focus on retaining voxels in regions with complex geometry and textures to better complement the rendering of the hybrid representation with the mesh.
-Inspired by PlenOctrees (Yu et al., 2021), we propose a voxel adjustment loss based on surface mesh weights.
+Inspired by PlenOctrees [6], we propose a voxel pruning loss based on surface mesh weights.
 Specifically, we calculate the rendering weight of mesh $w_{m}$ for rays hitting the surface and force it close to 1.
 The implementation is as follows:
-(7) $\mathcal{L}_{voxel}=\lambda_{voxel}\frac{1}{K}\sum_{k}(1-\text{exp}(w_{m}-1)),$
-where $\lambda_{voxel}$ represents the weight of the voxel adjustment loss, and $K$ is the number of samples taken in the rays hitting the surface.
+$\mathcal{L}_{voxel}=\lambda_{voxel}\frac{1}{K}\sum_{k}(1-\text{exp}(w_{m}-1)),$ (10)
+where $\lambda_{voxel}$ represents the weight of the voxel pruning loss, and $K$ is the number of samples taken before the rays hits the surface.
 The number of retained voxels is an important factor in determining the balance between rendering quality and speed.
 Concretely, retaining more voxels can enhance rendering quality but slow down rendering speed, while retaining fewer voxels can accelerate rendering speed but compromise rendering quality.
-Setting different voxel adjustment weights to limit the volumetric density during ray marching is one of the methods to achieve different trade-offs between speed and quality.
+Setting different voxel pruning weights to limit the volumetric density during ray marching is one of the methods to achieve different trade-offs between speed and quality.
 Additionally, voxels and mesh may overlap at the same spatial location, potentially affecting rendering quality and the volume of exported assets.
-To address this issue, we first calculate the occupancy grid of the meshes under the training poses.
-During voxel adjustment, we ignore the voxels in the mesh-occupied grid.
-We also adjust the ratio of sparse voxels in the hybrid representation by setting different resolutions $r_{mesh}$ for the mesh-occupancy grid.
-Table 1.
-Quantitative evaluation of the proposed method and SOTA methods with 3 outdoor scenes and 4 indoor scenes from Mip-NeRF 360 dataset (Barron et al., 2022).
-The methods that can render in real-time on mobile phones are highlighted.
-Outdoor Indoor GTX 1660Ti iPhone 14 Pro
-Methods Representations PSNR ↑ SSIM ↑ LPIPS ↓ PSNR ↑ SSIM ↑ LPIPS ↓ FPS @ 1920*1080 FPS @ 344*707
-NeRF MLP 21.46 0.458 0.515 26.84 0.790 0.370 - -
-Mip-NeRF 360 MLP 25.92 0.747 0.244 31.72 0.917 0.179 - -
-Instant-NGP Voxel 23.90 0.648 0.369 29.47 0.877 0.273 - -
-SNeRG++${}_{1024}$ Voxel 24.11 0.625 0.377 27.32 0.802 0.276 9.0 -
-MERF Voxel + Plane 24.41 0.678 0.298 27.80 0.855 0.271 9.7 25.3
-BakedSDF Mesh 23.40 0.577 0.351 27.20 0.845 0.300 131.0 -
-MobileNeRF Mesh 22.90 0.524 0.463 25.75 0.757 0.453 92.7 153.3
-NeRF2Mesh Mesh 22.74 0.497 0.463 25.20 0.721 0.340 231.0 210.0
-Vosh-Light Voxel + Mesh 23.63 0.586 0.392 27.02 0.792 0.294 79.7 70.0
-Vosh-Base Voxel + Mesh 23.74 0.592 0.389 27.07 0.791 0.291 38.0 33.7
-5. Real-Time Rendering
+To address this issue, we first calculate the occupancy grid of the meshes under the training poses and filter the voxels in the mesh-occupied grid out during rendering and baking.
+We also adjust the ratio of sparse voxels in the hybrid representation by setting different resolutions $r_{mesh}$ for the mesh-occupancy grid, which can also help suppress density of voxel component in Vosh.
+V Real-Time Rendering
 The real-time rendering based on Vosh is realized as a JavaScript 3D web application, leveraging the capabilities of the MERF real-time renderer.
 The rendering process is also hybrid and structured into two passes: the rasterization pass and the ray marching pass.
-Both passes are employed to make rendering result consistency between training on the GPU and inference on the web.
+Both passes are employed to make the rendering result consistent between training on the GPU and inference on the web.
 In the rasterization pass, meshes are uniformly shaded.
@@
 The output of the rasterization pass includes corresponding diffuse, view-dependent features and depth map.
-These output is then used in the subsequent ray marching pass.
-In the ray marching pass, we also employ a multi-resolution hierarchical structure for the occupancy grid introduced in SNeRG (Hedman et al., 2021) to mark the position of occupied voxels.
+This output is then used in the subsequent ray marching pass.
+In the ray marching pass, we also employ a multi-resolution hierarchical structure for the occupancy grid introduced in SNeRG [8] to mark the position of occupied voxels.
 This structure creates binary mask images in multiple resolution levels through max-pooling.
@@
 Leveraging the surface depth information from the rasterization pass proves instrumental in significantly diminishing the number of required sampling points during ray marching, thereby effectively reducing the computational cost associated with volume rendering.
-6. Experiments
+Figure 5: The rendering results and zoomed-in images in outdoor scenes obtained by our method, as well as some SOTA methods based on voxels or mesh representation.
+VI Experiments
 We have implemented real-time view synthesis using the proposed hybrid representation Vosh, and carried out comprehensive evaluations encompassing both rendering quality and speed.
-We also compare our method with state-of-the-art (SOTA) methods, like offline classical methods (NeRF (Mildenhall et al., 2021), Mip-NeRF 360 (Barron et al., 2022), Instant-NGP (Müller et al., 2022)) and real-time methods (Mobile-Nerf (Chen et al., 2023), NeRF2Mesh (Tang et al., 2023), BakedSDF (Yariv et al., 2023), MERF (Reiser et al., 2023), SNeRG++ (Hedman et al., 2021)).
-As done by SOTA methods, we used the challenging unbounded dataset Mip-NeRF 360 (Barron et al., 2022) for evaluation.
-Next, we elaborate the details of the experimental results.
-Figure 5. The rendering results and zoomed-in images obtained by our method, as well as some SOTA methods based on voxels or mesh representation.
-6.1. Rendering quality comparison
+We also compare our method with state-of-the-art (SOTA) methods, like the classical offline methods (NeRF [1], Mip-NeRF 360 [38], Instant-NGP [11]) and real-time methods (MobileNerf [13], NeRF2Mesh [14], BakedSDF [16], MERF [18], SNeRG++ [8]).
+As done by SOTA methods, we use the challenging unbounded dataset Mip-NeRF 360 [38] for evaluation.
+Here, we specifically list the two most relevant hybrid representation based methods of Adaptive Shells [19] and VMesh [20] as shown in Table I.
+For Adaptive Shells, we obtain quantitative results from its paper, while VMesh mostly handles individual objects or bounded scenes, while our method focuses more on real scenes.
+Next, we elaborate the details of the experiments.
+Figure 6: More rendering results and zoomed-in images in indoor scenes obtained by our method, as well as some SOTA methods based on voxels or mesh representation.
+VI-A Implementation details
+In the stage of training initial voxels, we use the classic MSE loss to measure pixel-level color differences.
+We use the same optimizer settings as MERF, e.g., 25000 steps for optimization and initially setting learning rates to 1e-2, which warmup on the first 100 steps and ultimately decay to 1e-3.
+We use 1024 as the resolution of the marching cubes and clean mesh with similar operations in NeRF2Mesh.
+In the stage of meshing density field, we use 20000 steps for mesh refinement.
+For the process of meshing density field, we set $p_{remove}=0.9$ for Vosh-Base and $p_{remove}=1.0$ for Vosh-Light.
+After the meshing stage, some messy floating faces may be obtained.
+We use similar parameters to repeat the operation of cleaning the mesh in the previous stage.
+And we encourage voxels to fill possible holes in the mesh by using L2 regularization terms with rendering weights for background colors.
+There are 20000 steps for Vosh optimization.
+For the Vosh-Base model, we set voxel pruning loss $\lambda_{voxel}=0.001$ and disable the binary mesh-occupancy grid.
+For the Vosh-Light model, we use voxel pruning loss $\lambda_{voxel}=0.01$ and the resolution $r_{mesh}=128$ for mesh occupancy grid as hyper-parameters.
+Besides, we did all experiments on a single NVIDIA RTX 4090.
+Three training stages take approximately 66, 35 and 58 minutes respectively.
+During the training process, the peak video memory usage for optimizing trainable parameters is about 11 Gigabytes.
+It should be noted though the training phase requires high-performance hardware, the inference phase can be performed directly on mobile devices with limited resources.
+VI-B Rendering quality comparison
 We conduct both quantitative and qualitative evaluation of rendering quality for view synthesis obtained by the proposed method and SOTA methods.
-Quantitative evaluation uses the metrics of PSNR, SSIM (Wang et al., 2004) and LPIPS (Zhang et al., 2018) for the assessment of rendering quality.
-Table 1 shows the statistics of the quantitative evaluation of different methods.
-Here, we classify these methods based on their ability to achieve real-time rendering on mobile devices like laptops and mobile phones.
-Methods such as SNeRG, MERF, and BakedSDF can only attain real-time rendering on select laptops with robust GPUs.
-It can be seen there exists an obvious gap between representations based on voxels and meshes.
-While BakedSDF narrows the quality gap with its millions of triangles, its substantial memory consumption makes it impractical for deployment on mobile phones.
+Quantitative evaluation uses the metrics of PSNR, SSIM [39] and LPIPS [40] for the assessment of rendering quality.
+Table I shows the statistics of the quantitative evaluation of different methods.
 Vosh demonstrates superior rendering quality in real-time rendering on mobile devices, even surpassing some methods that cannot achieve real-time rendering on mobile devices.
-For the outdoor scenes, Vosh enhances rendering quality over mesh-based methods by leveraging sparse voxels through the use of a contraction function (see Figure 5, 6 and 7).
-However, for the indoor scenes, methods have limitations in the representational capacity of tiny view-dependent MLPs.
+For the outdoor scenes, Vosh enhances rendering quality over mesh-based methods by leveraging sparse voxels through the use of a contraction function (see Figure 5 and 6).
+Rendering quality of Vosh in outdoor scenes also exceeds that of Adaptive Shells, which is also based on hybrid representation.
+The main reason is that Adaptive Shells constrain volume rendering to the regions near the surface, while in outdoor scenes, pure volume rendering is mostly of necessity.
+However, the limitations in representational capacity of the methods using tiny view-dependent MLPs limit rendering quality in indoor scenes.
 While Vosh achieves real-time rendering on mobile devices by employing a tiny view-dependent MLP, it has slightly lower rendering quality in indoor scenes compared to MLP-based methods.
-We provide a qualitative comparison with SOTA methods based on voxels (Reiser et al., 2023) and mesh (Tang et al., 2023) on the examples in Figure 5.
-We select one distant view and one close-up view for local magnification for better exhibition.
-The mesh-based methods often lack the ability to represent background regions far from the camera location, whereas voxel-based methods necessitate higher computational and storage costs to achieve enhanced rendering quality.
-Contrarily, our method improves rendering speed while essentially maintaining the rendering quality comparable to that of voxel-based methods.
-We refer the reader to the companion video for visual demonstrations of the results.
-Comparison with VMesh.
-It is worth noting that VMesh (Guo et al., 2023) is also built on the hybrid representation of volume and mesh, but primarily designed for rendering individual objects or scenes with boundaries, posing challenges in handling unbounded scenes like the examples in Figure 5 and 7.
-Nonetheless, our method produces better rendering results than VMesh, even for the synthetic bounded object as shown in Figure 6.
-Figure 6. The rendering results obtained by ours and VMesh (Guo et al., 2023).
-6.2. Rendering speed comparison
-We evaluated the performance of the methods applicable for rendering on mobile devices, including SNeRG++ (Hedman et al., 2021), MERF (Reiser et al., 2023), BakedSDF (Yariv et al., 2023), MobileNeRF (Chen et al., 2023), and NeRF2Mesh (Tang et al., 2023).
+We provide a qualitative comparison with SOTA methods based on voxels [18] and mesh [14] on the examples in Figure 5.
+We select one distant view and one close-up view with local magnification for better exhibition.
+The mesh-based methods often lack the ability to represent background regions far from the camera location, whereas volume-based methods necessitate higher computational and storage costs to achieve enhanced rendering quality.
+Contrarily, our method improves rendering speed while essentially maintaining the rendering quality comparable to that of volume-based methods.
+We refer the reader to the companion video for dynamic demonstrations of the results.
+VI-C Rendering speed comparison
+We evaluated the performance of the methods applicable for rendering on platform-independent mobile devices, including SNeRG++ [8], MERF [18], BakedSDF [16], MobileNeRF [13], and NeRF2Mesh [14].
+We conduct testing on iPhone 14 Pro equipped with the A16 chip with a resolution of 347$\times$600 and on Lenovo Yoga 14s 2021, which features Intel Iris Xe Graphics 80EU (28W, integrated in Intel Core i5-11350H) and operates with a resolution of 800$\times$600.
 For SNeRG++ and MERF, we follow their configurations with progressive upsampling turned off.
-We conduct testing on the iPhone 14 Pro equipped with the A16 chip with a resolution of 344*707 and on the Legion Y7000P 2019, which features NVIDIA GeForce GTX 1660Ti (Mobile version, 80W) and operates with a resolution of 1920*1080.
-Rendering speed tests for all comparative methods were conducted on the Mip-Nerf 360 dataset (Barron et al., 2022) in three outdoor scenes, using identical device settings, resolutions, and viewpoints.
-As shown on the right side of Table 1, our method can run with the real-time frame rate (¿30 FPS) on both iPhone 14 Pro and GTX 1660Ti.
-Particularly on the 1660Ti, our method demonstrates a considerable speed improvement compared to MERF and SNeRG.
-The latter two methods experience slow rendering speeds due to the necessity to sample each alive voxel traversed by every ray.
-BakedSDF is difficult to render on mobile phones and laptops with limited VRAM due to its substantial assets.
-MobileNeRF and NeRF2Mesh achieve rendering speeds that reach the maximum refresh rates of their respective devices (60Hz for Safari on iPhone 14 Pro and 144Hz for Legion Y7000P 2019).
+Rendering speed tests for all comparative methods are conducted on the Mip-Nerf 360 dataset [38] in three outdoor scenes, using identical device settings, resolutions, and viewpoints.
+As shown on the right side of Table I, our method can run with the real-time frame rate (>30 FPS) on both iPhone 14 Pro and Iris Xe 80EU.
+Adaptive Shell [19] can only achieve real-time rendering at 1080p resolution using an RTX 4090 GPU (36.18 FPS from its paper).
+Our method can achieve faster frame rates at the same resolution using an GTX 1660 Ti GPU.
+The volume-base methods have slow rendering speeds due to the necessity to sample each alive voxel traversed by every ray.
+BakedSDF [16] is difficult to render on mobile phones and laptops with limited VRAM due to its substantial assets.
+Specifically, BakedSDF has about 10 million triangles to represent unbounded scenes, mostly for complex backgrounds.
+Vosh, leveraging the efficiency of voxel component, reduces the need for triangles in low-quality and space-consuming background areas, using only about one-tenth as many.
+This greatly shortens the time of rasterization, providing an opportunity to enhance the performance based on sparse voxels.
+NeRF2Mesh [14] achieves rendering speeds that reach the maximum refresh rates of their respective devices (60Hz for Safari on iPhone 14 Pro and 90Hz for Lenovo Yoga 14s 2021).
 We make statistics on their potential rendering speeds using overlapping multiple passes.
-6.3. Ablation Study
-We conduct ablation experiments on the voxels to mesh conversion and voxel adjustment parts of the hybrid representation in Table 2.
+VI-D Comparison with VMesh
+It is worth noting that VMesh [20] is also built on the hybrid representation of volume and mesh, but primarily designed for rendering individual objects or scenes with boundaries, posing challenges in handling unbounded scenes like the examples in Figure 5.
+To highlight the effectiveness of the proposed method, we particularly compare Vosh with VMesh on NeRF-Synthetic dataset [1].
+The speed test in NeRF-Synthetic is based on the Lenovo Yoga 14s 2021 equipped with Iris Xe Graphics 80EU at a resolution of 800$\times$600.
+Due to the particularity of the NeRF-Synthetic dataset, the mesh extracted based on volume representation can usually achieve higher rendering quality than the volume representation itself after optimization [14].
+So we mesh the density field of the entire voxel grid, which means that the $p_{remove}$ is fixed to $1$.
+Our method has better performance than VMesh in the aspect of rendering quality or speed, as shown in Table II for quantitative comparison and Figure 7 for qualitative comparison.
+TABLE II: Quantitative evaluation of the proposed method and VMesh [20] in NeRF-Synthetic dataset [1].
+Methods PSNR SSIM LPIPS FPS
+VMesh 30.75 0.948 0.059 261.0
+Vosh-Base 30.96 0.948 0.064 85.0
+Vosh-Light 30.77 0.947 0.065 281.8
+Figure 7: The rendering results and zoomed-in images on the NeRF-Synthetic dataset [1] from our method and VMesh [20].
+Figure 8: The hybrid, mesh based and volume based rendering results obtained by our method and VMesh [20] on the NeRF-Synthetic dataset [1].
+The PSNR in the figure represents the evaluation result of the current images.
+We also provide the rendering results for the mesh part and the volume part separately, in addition to the rendering results of the hybrid representation of VMesh and Vosh, in Figure 8.
+Our reconstruction method of hybrid representation based on volumetric density field can restore more comprehensive scene information, thereby utilizing the extracted mesh to represent thin structures that are difficult to reconstruct in SDF, such as fishing nets and ropes in the figure.
+Hence, our method usually gains better rendering quality than VMesh.
+VI-E Storage comparison
+We conduct a statistical analysis of the graphics memory consumption and disk usage of the SOTA methods and the proposed method on the outdoor scenes of Mip-NeRF 360 [38], as shown in Table III.
+We use the occupying grid based approach proposed by SNeRG [8] to store voxels, and use methods similar to MobileNeRF [13] and NeRF2Mesh [14] to store mesh assets.
+Since we do not use the triplane representation which is the core contribution of MERF [18], but only recover the scene information and generate a hybrid representation based on the classic voxel representation SNeRG [8], it appears to be slightly larger than MERF in terms of disk storage.
+Although our method does not make any additional optimizations in terms of memory, it still has certain advantages compared to methods with similar rendering quality or speed.
+TABLE III: Memory consumption and disk usage of the SOTA methods in outdoor scenes on the Mip-NeRF 360 dataset.
+Methods VRAM (MB) ↓ Disk (MB) ↓
+SNeRG++ 340 624
+MERF 360 149
+NeRF2Mesh 426 123
+MobileNeRF 927 345
+BakedSDF 953 807
+3DGS 3701 1270
+Vosh-Light 490 232
+Vosh-Base 506 488
+VI-F Ablation study
+We conduct ablation experiments on the meshing density field and voxel pruning parts of the hybrid representation on outdoor scenes of Mip-NeRF dataset [38], as shown in Table IV and Table V.
 The assessment involves a quantitative comparison, considering both rendering quality and speed.
-The ‘full bound’ row refers to the results of the variant that do not use coarse selection of voxels in the voxels to mesh conversion.
-The experimental results show that converting all voxels containing surfaces to mesh not only reduces the rendering quality of the final hybrid representation, but also significantly increases the storage of the assets.
-In the voxel adjustment stage, we compare the complete model with variants of Vosh-Base, incorporating different loss weights or resolutions for mesh occupancy grids.
-The rows with $\lambda_{voxel}$ show the results of different hyper-parameter weights of the loss function for voxel adjustment.
-The rows with $r_{mesh}$ indicate the results of using mesh occupancy grids with different resolutions.
-For the Vosh-Base model, we use $\lambda=0.001$ and $r_{mesh}=128$ as hyper-parameters.
-For the Vosh-Light model, we use $\lambda=0.1$ and $r_{mesh}=32$ as hyper-parameters.
-Experimental results suggest that increasing the weight of voxel adjustment loss or reducing the resolution of the mesh occupancy grid aids in compressing the number of voxels, leading to speed benefits at the expense of rendering quality.
-Table 2. The ablation study in the construction of hybrid representation.
-Variants PSNR ↑ SSIM ↑ LPIPS ↓ FPS ↑
-Vosh-Base 23.74 0.592 0.389 33.7
-full bound 23.49 0.583 0.395 36.0
-$\lambda_{voxel}=0$ 23.73 0.591 0.389 30.0
-$\lambda_{voxel}=0.01$ 23.72 0.590 0.389 38.7
-$\lambda_{voxel}=0.1$ 23.69 0.590 0.390 46.7
-$r_{mesh}=0$ 23.80 0.593 0.386 32.0
-$r_{mesh}=1024$ 23.79 0.593 0.387 32.0
-$r_{mesh}=64$ 23.70 0.590 0.389 37.7
-Vosh-Light 23.63 0.586 0.392 70.0
-6.4. Limitations
+We compare the variants of different optimization methods for differentiable surface optimization in Table IV.
+In addition to rendering quality, we also report the number of faces (in millions) of different variants .
+It can be seen that optimizing the geometry of the mesh based on differentiable rendering can significantly improve rendering quality.
+The experiments on surface operations indicate that it can better preserve rendering quality without mesh simplification, but might result in a large number of triangles.
+It might also cause some loss of rendering quality without performing mesh subdivision.
+So we use mesh subdivision and simplification based on rendering quality and normal angle change to significantly improve rendering quality while decreasing the number of triangles.
+Compared with the mesh optimization in NeRF2Mesh [14], our method additionally considers the angle of normals between adjacent triangles during the optimization, which can improve rendering quality while reducing the number of triangles.
+At the same time, our meshing strategy based on rendering quality at the same stage also helps generating the high quality hybrid representation as shown in Fig 9.
+TABLE IV: The ablation study in the differentiable surface refinement.
+Variants PSNR SSIM LPIPS # faces
+Mesh Baseline 22.04 0.475 0.448 1.76
+Refine w/o subd 23.52 0.597 0.379 1.44
+Refine w/o deci 23.61 0.601 0.377 2.14
+Refine w/o norm 23.54 0.594 0.381 1.83
+Refine 23.57 0.598 0.379 1.69
+TABLE V: The ablation study in the construction of hybrid representation.
+Variants PSNR SSIM LPIPS # faces # voxels
+Meshing $p_{remove}=0.5$ 24.49 0.642 0.357 0.81 98999
+$p_{remove}=1.0$ 23.95 0.600 0.384 1.69 47631
+Refined mesh 23.57 0.598 0.379 1.69 0
+Voxel grid 24.48 0.646 0.358 0 126533
+Voxel prune. $\lambda_{voxel}=0$ 24.23 0.625 0.358 1.41 74317
+$\lambda_{voxel}=0.01$ 23.90 0.606 0.377 1.41 36255
+$r_{mesh}=256$ 23.80 0.604 0.377 1.41 41743
+$r_{mesh}=512$ 23.94 0.609 0.373 1.41 52280
+Vosh-Base 24.18 0.623 0.364 1.41 57438
+Vosh-Light 23.74 0.602 0.377 1.41 2489
+Figure 9: The hybrid, mesh based and volume based rendering results obtained by our method on Mip-NeRF 360 dataset [38]. The PSNR in the figure represents the evaluation result of the current images.
+To better illustrate the effectiveness of each component in the construction process of our hybrid representation, we compare Vosh-Base with its variants in Table V, which includes the threshold $p_{remove}$ in meshing based on rendering quality, the weight of the voxel pruning loss $\lambda_{voxel}$, and the resolution $r_{mesh}$ for mesh occupancy grid.
+We also report the number of faces (in millions) that make up the mesh component and the number of alive voxels that make up the voxels component in the Vosh representation.
+In the section of meshing density field, we compare the variants which $p_{remove}=0.5$ and $1.0$ with the Vosh-Base model with $p_{remove}=0.9$.
+Meanwhile, in addition to the rendering results of the initial voxel grid, we also provide refined rendering results for all surfaces obtained by meshing the entire voxel grid.
+When setting $p_{remove}=1.0$, it means that the entire density field is meshed, not based on rendering quality.
+The results show that this can only achieve a small improvement in quality.
+When setting $p_{remove}=0.5$, our hybrid representation that combines the advantages of voxel and mesh components achieve higher or comparable quality rendering results than the initial voxel grid.
+In the section of voxel pruning, we compare the variants $\lambda_{voxel}=0$ and $0.01$ with the Vosh-Base model with $\lambda_{voxel}=0.001$.
+We also list the rendering results of the variants $r_{mesh}=256$ and $512$, while Vosh-Base model turns off the check of mesh occupancy grid.
+Experimental results suggest that increasing the weight of voxel pruning loss or reducing the resolution of the mesh occupancy grid aids in compressing the number of alive voxels, leading to speed benefits at the expense of rendering quality.
+At the same time, we also provide the hybrid, mesh based and volume based rendering results of several representative variants of Vosh, as shown in Figure 9.
+We use Vosh-Base in the second line as the benchmark, which sets $p_{remove}$ to 0.9 and $\lambda_{voxel}$ to 0.001.
+As a comparison, the first line only changes $p_{remove}$ to 0.5 and the third line only changes $\lambda_{voxel}$ to 0.01.
+On the basis of Vosh-Base, Vosh-Light sets $\lambda_{voxel}$ to 0.01, $p_{remove}$ to 1.0 and $r_{mesh}$ to 128.
+Based on the qualitative comparison in this figure, it can be seen that we can extract surfaces with higher reconstruction quality by determining the region where the density field meshes based on the difference in rendering quality through the threshold $p_{remove}$.
+And the balance between quality and speed in the hybrid representation model can be adjusted by setting the different weight $\lambda_{voxel}$ of the voxel pruning loss and the resolution $r_{mesh}$ for mesh occupancy grid.
+VI-G Limitations
 Although the proposed hybrid representation Vosh brings valuable advantages to real-time view synthesis, it is not without limitations.
@@
 These limitations, in turn, may result in potential degradation of mesh optimization quality, thereby impacting the rendering quality based on the hybrid representation.
-7. Conclusion
+VII Conclusion
 We have presented a hybrid representation Vosh, composed of both voxels and mesh, for NeRF-based real-time view synthesis.